# Bi-directional Deformation for Parameterization of Neural Implicit Surfaces

### v1

@@ Abstract
The growing capabilities of neural rendering have increased the demand for new techniques that enable the intuitive editing of 3D objects, particularly when they are represented as neural implicit surfaces. In this paper, we present a novel neural algorithm to parameterize neural implicit surfaces to simple parametric domains, such as spheres, cubes or polycubes, where 3D radiance field can be represented as a 2D field, thereby facilitating visualization and various editing tasks. Technically, our method computes a bi-directional deformation between 3D objects and their chosen parametric domains, eliminating the need for any prior information. We adopt a forward mapping of points on the zero level set of the 3D object to a parametric domain, followed by a backward mapping through inverse deformation. To ensure the map is bijective, we employ a cycle loss while optimizing the smoothness of both deformations. Additionally, we leverage a Laplacian regularizer to effectively control angle distortion and offer the flexibility to choose from a range of parametric domains for managing area distortion. Designed for compatibility, our framework integrates seamlessly with existing neural rendering pipelines, taking multi-view images as input to reconstruct 3D geometry and compute the corresponding texture map. We also introduce a simple yet effective technique for intrinsic radiance decomposition, facilitating both view-independent material editing and view-dependent shading editing. Our method allows for the immediate rendering of edited textures through volume rendering, without the need for network re-training. Moreover, our approach supports the co-parameterization of multiple objects and enables texture transfer between them. We demonstrate the effectiveness of our method on images of human heads and man-made objects. We will make the source code publicly available.
@@ 1 Introduction
Neural radiance fields (NeRF)(Mildenhall et al., 2020) have garnered remarkable success in both the computer vision and computer graphics communities, redefining benchmarks for high-quality renderings and novel view synthesis. Building upon NeRF, a variety of methods leveraging implicit neural representations (Oechsle et al., 2021; Wang et al., 2021; Yariv et al., 2021) have emerged, delivering high-fidelity 3D reconstructions. The demand for intuitive 3D object editing techniques has increased with the growing capabilities of neural rendering, especially for neural implicit surfaces. Extensive explorations have been made in editing within the Neural Rendering framework to achieve diverse visual effects. We can implicitly alter the shape and color of 3D objects by manipulating the latent codes of the network(Liu et al., 2021; Yenamandra et al., 2021; Xu et al., 2023). In the work of (Tojo & Umetani, 2022; Kuang et al., 2023), palette-based methods have been employed to facilitate transformations in the color and style of rendered scenes. However, these methods suffer from limited controllability and disallow local pixel editing. Seal3D (Wang et al., 2023) proposed a teacher-student training strategy to edit the interactive scene at the pixel level. Yet, this method is tricky to handle shading, and each subsequent edit required a new training iteration. By employing neural networks to formulate a differentiable surface parameterization, it can be facilitated to edit textures or shapes within the neural rendering pipeline (Xiang et al., 2021; Ma et al., 2022). Existing neural parameterization techniques still face significant challenges, such as the need for appropriate distortion constraints to ensure accurate detail reconstruction or reducing the reliance on prior information from tracked mesh and UV mapping. The final quality of edited visual effects is significantly impacted by these issues.
@@
This paper aims at developing an intuitive and easy-to-use tool for appearance editing in neural implicit surfaces. Towards this goal, we propose a novel method to parameterize neural implicit surfaces to parametric domains, such as spheres, cubes or polycubes, where the radiance field is represented as a texture map. Technically, our parameterization algorithm utilizes a neural network to learn a bi-directional deformation between 3D objects and the chosen parametric domains. This involves a forward deformation that maps points from the zero level set of the neural implicit surface to the parametric domain, followed by an inverse deformation, mapping points backward. Notably, we do not require any explicit prior information in learning both deformations. We employ a cycle loss to ensure the smoothness of the bi-directional deformation, and a Laplacian regularization to effectively control angle distortion. Our method also supports the flexible selection of parametric domains for controlling area distortion. With the parametrization, 3D radiance field becomes essentially a 2D field, facilitating visualization and various editing tasks. we further decompose the 2D radiance field into two components: view-independent material and view-dependent shading, streamlining both texture and shading editing. Our neural parameterization algorithm is fully compatible with existing neural rendering pipelines, allowing 3D reconstruction from multi-view images as input and the creation of texture maps simultaneously. It allows for the immediate rendering of edited textures through volume rendering, without the need for network re-training. Moreover, it also supports co-parameterization of multiple objects of similar geometry and enable texture transfer between them. We validate the effectiveness of our method on images of human heads and man-made objects.
Our contributions are summarized as follows:
@@
  • We present a neural parameterization framework that computes bi-directional deformation between 3D objects and their parametric domains, eliminating the need for any prior information from tracked mesh or UV mapping. Our approach utilizes Laplacian loss to minimize angle distortion and provides a choice of parametric domains, such as spheres, cubes, or polycubes, to control area distortion.
@@
  • Our framework is fully compatible with existing neural rendering architectures, accepting multi-view images as input to reconstruct 3D geometry and generate UV maps as output. Additionally, it enables the direct rendering of modified textures using volume rendering pipeline.
@@
Figure 1: Our method takes multi-view images as input to reconstruct 3D geometry as neural implicit surfaces, while representing appearance through a mapping between the 3D surface and a user-selected parametric domain. Such a representation allows volume rendering of modified textures without the need of network retraining, facilitating both view-dependent shading editing and view-independent material editing.
@@ 2 Related Work
Parameterization. Surface parameterization (Sheffer et al., 2007; Floater & Hormann, 2005) aims at computing a bijective mapping between the 3D surface and a suitable parametric domain, usually a 2D region or a simple 3D object, such as spheres (Gotsman et al., 2003) and polycubes (García et al., 2013). Serving as an important computational tool in computer graphics and digital geometry processing (Sheffer et al., 2007), surface parameterization facilitates various applications, including texture editing (Fang & Hart, 2004), surface painting (Sun et al., 2013), details transfer (Biermann et al., 2002), remeshing (Praun & Hoppe, 2003), among others. Neural networks are now commonly used in digital geometry processing, resulting in the creation of parameterization algorithms based on deep learning techniques (Groueix et al., 2018; Williams et al., 2019; Bednarik et al., 2020; Guo et al., 2022). Unlike classical methods that yield global and seamless parametrization, most neural parameterization methods are computing local parametrization (Groueix et al., 2018; Zhang et al., 2023; 2022). These methods typically partition the 3D model into multiple patches and parameterize each to a 2D region, possibly followed by a post-processing step to stitch the patches together. Although these methods are efficient and can handle surfaces of arbitrary geometry and topology, they frequently encounter problems including non-bijectivity, a lack of smoothness, the presence of seams and overlaps, and large distortions. There are also works on neural parametric surfaces, defined in rectangular domains (Low & Lee, 2022) and $n$-sided patches (Yang et al., 2023). They can model complex surface geometries with high precision; however, such representations are not applicable to neural implicit surfaces and thereby incompatible with neural rendering.
@@
Neural Implicit Functions. Recent years have seen a surge in the successful application of neural implicit functions within the realm of 3D deep learning. Compared to explicit representations such as point clouds (Qi et al., 2017), voxels (Qi et al., 2016), and meshes (Wang et al., 2018), neural implicit representations offer unique advantages including flexibility, continuity, and robustness (Mescheder et al., 2019; Saito et al., 2019; Park et al., 2019). When integrated within the neural rendering pipeline, neural implicit surfaces exhibit exceptional quality in 3D reconstructions from multi-view images (Li & Zhang, 2021; Wang et al., 2022a; b; Xu et al., 2023; Müller et al., 2022; Li et al., 2023; Rosu & Behnke, 2023). However, as geometries and radiance fields are encoded as network parameters, editing them becomes complex and non-intuitive, compared to explicit representations. This paper aims at tackling this challenge by explicitly representing radiance fields as texture mapping that encompasses both view-dependent and view-independent components.
@@
NeRF Editing. Numerous studies have focused on editing neural radiance fields, including relighting (Srinivasan et al., 2021; Li et al., 2022), composition (Pérez et al., 2023; Yang et al., 2021), content generation (Niemeyer & Geiger, 2021), and shape editing (Yuan et al., 2023; Lin et al., 2022). Editing tools operating at the scene- or object-level, such as (Yang et al., 2021; Yenamandra et al., 2021; Huang et al., 2022), utilize latent codes to modify and stylize the appearance of an object or entire scenes. Conversely, pixel-level editing tools, such as Seal3D (Wang et al., 2023) and NeuMesh (Yang et al., 2022), utilize training-based approaches to edit fine-grained details. However, they necessitate retraining for each editing, lacking efficiency and robustness. Parameterization-based methods, such as NeuTex (Xiang et al., 2021), ISO-UVField (Sagnik Das & Samaras, 2022) and NeP (Ma et al., 2022), integrate the learning of UV mapping into the neural rendering framework, providing an intuitive means for editing within 2D domains. However, these methods often suffer from large distortions in the parametrization. Furthermore, they require 3D prior information from tracked meshes and UV mapping - conditions that may not be readily met in practical, real-world scenarios. IntrinsicNeRF (Ye et al., 2023) enhances standard neural radiance fields by generating additional outputs, including reflectance, shading, and a residual term. Incorporating a semantic branch, it facilitates real-time scene editing. However, it often results in aggregated colors, akin to palette-based methods (Tojo & Umetani, 2022; Kuang et al., 2023), hence cannot offer fine-grained editing results at the pixel level. Our approach decomposes the radiance field into view-dependent shadings and view-independent materials. Both components are represented as textures, enabling intuitive material modifications while preserving consistent shading.
@@ 3 Method
[…]
The input of our method is a collection of RGB images, denoted =$\mathcal{I}=\{\mathbf{I}_{i}\}$, representing either multiple objects of similar geometry or a single object, captured from different viewpoints. To differentiate among objects, we assign a latent shape code $\mathbf{z}_{s}$ and an appearance code $\mathbf{z}_{a}$ to each object. This facilitates the co-parameterization of multiple objects and enables transfer material and shading among them. As illustrated in Figure 2, our method reconstructs 3D geometry as a neural implicit surface and encodes the 3D radiance field in a simple parametric domain, such as spheres and polycubes. For each input image, we randomly select a set of pixels. Then, for each pixel, we shoot a ray originating from the location of camera and traversing through the pixel. If the ray intersects with the surface, at a point, denoted as $p$, we compute its corresponding position in the parametric domain $\mathcal{D}$, denoted as $\mathbf{p}^{\prime}$, via a bi-directional deformation (Sec. 3.1): the forward deformation (depicted in orange) generates a displacement vector so that the updated position $\mathbf{p}^{\prime}$ is in the parametric domain; subsequently, the inverse deformation (depicted in green) maps $\mathbf{p}^{\prime}$ back to a point $\mathbf{p}^{\prime\prime}$, which is anticipated to be in close proximity to the surface. We incorporate a cycle loss to penalize occurrences where $\mathbf{p}^{\prime\prime}$ deviates from $\mathbf{p}$. In addition, we employ a Laplacian loss (Sec. 3.2) to effectively reduce the angle distortion. Furthermore, we decompose the radiance into a view-independent material field and a view-dependent shading field (Sec. 3.3), each of which can be edited independently, thus augmenting the editing capability of our framework.
@@ 3.1 Bi-directional deformation
The bi-directional deformation module serves as the pivotal component within our neural parameterization framework. As illustrated in Figure 2, this module consists of two sub-networks: $F_{\text{def}}$, which is responsible for the forward mapping from the original model to the parameter domain, and $F_{\text{inv-def}}$, which establishes the inverse mapping. Let $\mathcal{S}$ be the 3D surface and $\mathcal{D}$ be the parametric domain; both are represented by the zero level-set of neural signed distance fields. For an arbitrary point $\mathbf{p}\in\mathcal{S}$, the forward deformation $F_{\text{def}}(\mathbf{p},\mathbf{z}_{s})$ computes a displacement vector, mapping the point $\mathbf{p}$ to a corresponding point $\mathbf{p}^{\prime}\in\mathcal{D}$ in the parametric domain as follows:
@@
In training the networks $F_{\text{def}}$ and $F_{\text{inv-def}}$, our objective is to achieve a bijective mapping between the 3D surface $\mathcal{S}$ and the parametric domain $\mathcal{D}$. To achieve this goal, we adopt a cycle loss, which is similar to NeuTex (Xiang et al., 2021):
$\mathcal{L}_{\text{cycle}}=\sum_{\mathbf{p}\in\mathcal{S}}\lambda(\mathbf{p})\|\mathbf{p}-\mathbf{p}^{\prime\prime}\|_{2},$ (4)
where the weight $\lambda$ characterizes the importance of sample $\bf p$ to the loss. The reason that we do not consider all sample points equally is that a sample point may not be exactly on the zero level-set of the SDF. To tolerate such an inaccuracy, we define $\lambda(\mathbf{p})=T(\mathbf{p})(1-\exp(-\sigma(s)\delta(\textbf{p})))$ as the color weight used in volume rendering (Mildenhall et al., 2020). Here, $T(\mathbf{p})$ is the transparency of sample $\mathbf{p}$ in the viewing direction, $\delta$ is the length of the sample interval along the ray, and $s$ is the signed distance value of the point $\mathbf{p}$. Clearly, when $\bf p$ is away from the surface, the coefficient $\lambda(\mathbf{p})$ has few effects on this loss. We refer the readers to (Mildenhall et al., 2020) for details about volume rendering and the computation of transparency $T$. By explicitly minimizing the distance between $\mathbf{p}$ and $\mathbf{p}^{\prime\prime}$, the cycle loss effectively prevents scenarios where two distinct points $\mathbf{p}_{1},\mathbf{p}_{2}\in\mathcal{S}$ map to an identical point $\mathbf{p}^{\prime}\in\mathcal{D}$. In such cases, the inverse deformation $F_{\text{inv-def}}$ would be unable to map the single point $\mathbf{p}^{\prime}$ back to the respective distinct points $\mathbf{p}_{1}^{\prime\prime}$ and $\mathbf{p}_{2}^{\prime\prime}$.

### v2

# Parameterization-driven Neural Implicit Surfaces Editing

@@ Abstract
The growing capabilities of neural rendering have increased the demand for new techniques that enable intuitive editing of 3D objects, particularly when they are represented as neural implicit surfaces. In this paper, we present a novel neural algorithm to parameterize neural implicit surfaces to simple parametric domains, such as spheres, cubes, or polycubes, thereby facilitating visualization and various editing tasks. Technically, our method computes a bi-directional deformation between 3D objects and their chosen parametric domains, eliminating the need for any prior information. We adopt a forward mapping of points on the zero level set of the 3D object to a parametric domain, followed by a backward mapping through inverse deformation. To ensure the map is bijective, we employ a cycle loss while optimizing the smoothness of both deformations. Additionally, we leverage a Laplacian regularizer to effectively control angle distortion and offer the flexibility to choose from a range of parametric domains for managing area distortion. Designed for compatibility, our framework integrates seamlessly with existing neural rendering pipelines, taking multi-view images as input to reconstruct 3D geometry and compute the corresponding texture map. We also introduce a simple yet effective technique for intrinsic radiance decomposition, facilitating both view-independent material editing and view-dependent shading editing. Our method allows for the immediate rendering of edited textures through volume rendering, without the need for network re-training. Moreover, our approach supports the co-parameterization of multiple objects and enables texture transfer between them. We demonstrate the effectiveness of our method on images of human heads and man-made objects. We will make the source code publicly available.
Figure 1: Overview of NeRF-based editing work. We summarize the existing work into three main categories: manipulating the latent codes, re-training, and parameterization. In this study, we propose a novel NeRF editing method based on our 3D neural parameterization framework.
@@ 1 Introduction
Neural radiance fields (NeRF) [29] have garnered remarkable success in both the computer vision and computer graphics communities, redefining benchmarks for high-quality renderings and novel view synthesis. Building upon NeRF, a variety of methods leveraging implicit neural representations [32, 47, 61] have emerged, delivering high-fidelity 3D reconstructions. The demand for intuitive 3D object editing techniques has increased with the growing capabilities of neural rendering, especially for neural implicit surfaces. Extensive explorations have been made in editing within the Neural Rendering framework to achieve diverse visual effects. By manipulating the latent codes of the network, one can implicitly alter the shape and color of 3D objects[25, 63, 54]. In the work of [45, 20], palette-based methods have been employed to facilitate transformations in the color and style of rendered scenes. However, these methods suffer from limited controllability and disallow local pixel editing. Seal3D [48] proposed a teacher-student training strategy to edit the interactive scene at the pixel level. Yet, this method is tricky to handle shading, and each subsequent edit required a new training iteration. By employing neural networks to formulate a differentiable surface parameterization, it can be facilitated to edit textures or shapes within the neural rendering pipeline [53, 27]. Existing neural parameterization techniques still face significant challenges, such as the need for appropriate distortion constraints to ensure accurate detail reconstruction or reducing the reliance on prior information from tracked mesh and UV mapping. The final quality of edited visual effects is significantly impacted by these issues. We summarize the main work related to NeRF-editing in Fig. 1.
@@
This paper aims at developing an intuitive and easy-to-use tool for appearance editing in neural implicit surfaces. Towards this goal, we propose a novel method to parameterize neural implicit surfaces to parametric domains where the radiance field is represented as a texture map. Distinguished from previous work that using 2D texture mapping, our neural parameterization algorithm supports the selection of different 3D parametric domains according to geometric shapes to avoid seams and reduce distortion, such as spheres, cubes, and polycubes. Technically, our parameterization algorithm utilizes a neural network to learn a bi-directional deformation between 3D objects and the chosen parametric domains. This involves a forward deformation that maps points from the zero level set of the neural implicit surface to the parametric domain, followed by an inverse deformation, mapping points backward. Notably, we do not require any explicit prior information in learning both deformations. We employ a cycle loss to ensure the smoothness of the bi-directional deformation, and a Laplacian regularization to effectively control angle distortion. Our method also supports the flexible selection of parametric domains for controlling area distortion. With the parametrization, 3D radiance field is mapped to a simple 3D parametric domain which can be further disassembled into a 2D domain, facilitating visualization and various editing tasks. Furthermore, we decompose the 2D radiance field into two components: view-independent material and view-dependent shading, streamlining both texture and shading editing. Our neural parameterization algorithm is fully compatible with existing neural rendering pipelines, allowing 3D reconstruction from multi-view images as input and the creation of texture maps simultaneously. It allows for the immediate rendering of edited textures through volume rendering, without network re-training. Moreover, it supports co-parameterization of objects of similar geometry and enables texture transfer between them. We validate the effectiveness of our method on scenes of human heads and man-made objects. Our contributions are summarized as follows:
@@
  • We present a neural parameterization framework that computes bi-directional deformation between 3D objects and their parametric domains, eliminating the need for any prior information from tracked mesh or UV mapping. Our approach utilizes Laplacian loss to minimize angle distortion and provides a choice of 3D parametric domains, such as spheres, cubes, or polycubes, to control area distortion.
@@
  • Our neural parameterization framework is fully compatible with existing neural rendering architectures, accepting multi-view images as input to reconstruct 3D geometry and generate UV maps as output. Additionally, it enables the direct rendering of modified textures using the volume rendering pipeline.
@@
[…]
@@ 2 Related Work
Parameterization. Surface parameterization [41, 6] aims at computing a bijective mapping between the 3D surface and a suitable parametric domain, usually a 2D region or a simple 3D object, such as spheres [9] and polycubes [8, 44]. Serving as an important computational tool in computer graphics and digital geometry processing [41], surface parameterization facilitates various applications, including texture editing [5], surface painting [43], details transfer [2], remeshing [35], among others. Neural networks are now commonly used in digital geometry processing, resulting in the creation of parameterization algorithms based on deep learning techniques [11, 51, 1, 14]. Unlike classical methods that yield global and seamless parametrization, most neural parameterization methods are computing local parametrization [11, 67, 66]. These methods typically partition the 3D model into multiple patches and parameterize each to a 2D region, possibly followed by a post-processing step to stitch the patches together. Although these methods are efficient and can handle surfaces of arbitrary geometry and topology, they frequently encounter problems including non-bijectivity, a lack of smoothness, the presence of seams and overlaps, and large distortions. There are also works on neural parametric surfaces, defined in rectangular domains [26] and $n$-sided patches [59]. They can model complex surface geometries with high precision; however, such representations are not applicable to neural implicit surfaces and thereby incompatible with neural rendering.
@@
Neural Implicit Functions. Recent years have seen a surge in the successful application of neural implicit functions within the realm of 3D deep learning. Compared to explicit representations such as point clouds [37], voxels [36], and meshes [46], neural implicit representations offer unique advantages including flexibility, continuity, and robustness [28, 40, 33]. When integrated within the neural rendering pipeline, neural implicit surfaces exhibit exceptional quality in 3D reconstructions from multi-view images [21, 49, 50, 54, 30, 22, 38]. However, as geometries and radiance fields are encoded as network parameters, editing them becomes complex and non-intuitive, compared to explicit representations. This paper aims at tackling this challenge by explicitly representing radiance fields as texture mapping that encompasses both view-dependent and view-independent components.
@@
NeRF Editing. Numerous studies have focused on editing neural radiance fields, including relighting [42, 23], composition [34, 56], content generation [31], and shape editing [65, 24]. Editing tools operating at the scene- or object-level, such as [56, 63, 16], utilize latent codes to modify and stylize the appearance of an object or entire scenes. Conversely, pixel-level editing tools, such as NeuMesh [57] and Seal3D [48], utilize training-based approaches to edit fine-grained details. However, NeuMesh relies on explicit meshes which lack efficiency and robustness. Although Seal3D achieves interactive editing, it still suffers from poor geometry reconstruction and cannot change the illumination scene. Parameterization-based methods, such as NeuTex [53], ISO-UVField [39] and NeP [27], integrate the learning of UV mapping into the neural rendering framework, providing an intuitive means for editing within 2D domains. However, these methods often suffer from large distortions in the parametrization. Furthermore, they require 3D prior information from tracked meshes and UV mapping - conditions that may not be readily met in practical, real-world scenarios. IntrinsicNeRF [62] enhances standard neural radiance fields by generating additional outputs, including reflectance, shading, and a residual term. Incorporating a semantic branch, it facilitates real-time scene editing. However, it often results in aggregated colors, akin to palette-based methods [45, 20], hence cannot offer fine-grained editing results at the pixel level. Our approach decomposes the radiance field into view-dependent shadings and view-independent materials. Both components are represented as textures, enabling intuitive material modifications while preserving consistent shading.
@@ 3 Method
The input of our method is a collection of RGB images, denoted =$\mathcal{I}=\{\mathbf{I}_{i}\}$, representing either multiple objects of similar geometry or a single object, captured from different viewpoints. To differentiate among objects, we assign a latent shape code $\mathbf{z}_{s}$ and an appearance code $\mathbf{z}_{a}$ to each object. This facilitates the co-parameterization of multiple objects and enables the transfer of material and shading among them. As illustrated in Figure 2, our method reconstructs 3D geometry as a neural implicit surface and encodes the 3D radiance field in a simple parametric domain, such as spheres and polycubes. For each input image, we randomly select a set of pixels. Then, for each pixel, we shoot a ray originating from the location of camera and traversing through the pixel. If the ray intersects with the surface, at a point, denoted as $\mathbf{p}$, we compute its corresponding position in the parametric domain $\mathcal{D}$, denoted as $\mathbf{p}^{\prime}$, via a bi-directional deformation (Sec. 3.1): the forward deformation (depicted in orange) generates a displacement vector so that the updated position $\mathbf{p}^{\prime}$ is in the parametric domain; subsequently, the inverse deformation (depicted in green) maps $\mathbf{p}^{\prime}$ back to a point $\mathbf{p}^{\prime\prime}$, which is anticipated to be in close proximity to the surface. We incorporate a cycle loss to penalize occurrences where $\mathbf{p}^{\prime\prime}$ deviates from $\mathbf{p}$. In addition, we employ a Laplacian loss (Sec. 3.2) to effectively reduce the angle distortion between the parametric domain and the original surface. Furthermore, we decompose the radiance into a view-independent material field and a view-dependent shading field (Sec. 3.3), each of which can be edited independently, thus augmenting the editing capability of our framework.
GT Normal $\mathcal{D}$ $\mathcal{S}$
Figure 3: Parameterization results. Human heads are co-parameterized to a sphere owing to the clear geometric resemblance (top four rows). Given the diverse geometry of cars, each is parameterized to a polycube domain (bottom three rows). From left to right: the input image, the normal map of the reconstructed geometry $\mathcal{S}$, the parametric domain $\mathcal{D}$, and the reconstructed surface $\mathcal{S}$. See supplementary material for additional results.
@@ 3.1 Bi-directional deformation
The bi-directional deformation module serves as the pivotal component within our neural parameterization framework. As illustrated in Figure 2, this module consists of two sub-networks: $F_{\text{def}}$, which is responsible for the forward mapping from the original model to the parameter domain, and $F_{\text{inv-def}}$, which establishes the inverse mapping.
Let $\mathcal{S}$ be the 3D surface and $\mathcal{D}$ be the parametric domain; both are represented by the zero level-set of neural signed distance fields. For an arbitrary point $\mathbf{p}\in\mathcal{S}$, the forward deformation $F_{\text{def}}(\mathbf{p},\mathbf{z}_{s})$ computes a displacement vector, mapping the point $\mathbf{p}$ to a corresponding point $\mathbf{p}^{\prime}\in\mathcal{D}$ in the parametric domain as follows:
@@
In training the networks $F_{\text{def}}$ and $F_{\text{inv-def}}$, our objective is to achieve a bijective mapping between the 3D surface $\mathcal{S}$ and the parametric domain $\mathcal{D}$. To achieve this goal, we adopt a cycle loss, which is similar to NeuTex [53]:
$\mathcal{L}_{\text{cycle}}=\sum\nolimits_{\mathbf{p}\in\mathcal{S}}\lambda(\mathbf{p})\|\mathbf{p}-\mathbf{p}^{\prime\prime}\|_{2},$ (4)
where the weight $\lambda$ characterizes the importance of sample $\bf p$ to the loss. The reason that we do not consider all sample points equally is that a sample point may not be exactly on the zero level-set of the SDF. To tolerate such an inaccuracy, we define $\lambda(\mathbf{p})=T(\mathbf{p})(1-\exp(-\sigma(s)\delta(\textbf{p})))$ as the color weight used in volume rendering [29]. Here, $T(\mathbf{p})$ is the transparency of sample $\mathbf{p}$ in the viewing direction, $\delta$ is the length of the sample interval along the ray, and $s$ is the signed distance value of the point $\mathbf{p}$. Clearly, when $\bf p$ is away from the surface, the coefficient $\lambda(\mathbf{p})$ has few effects on this loss. We refer the readers to [29] for details about volume rendering and the computation of transparency $T$.
By explicitly minimizing the distance between $\mathbf{p}$ and $\mathbf{p}^{\prime\prime}$, the cycle loss effectively prevents scenarios where two distinct points $\mathbf{p}_{1},\mathbf{p}_{2}\in\mathcal{S}$ map to an identical point $\mathbf{p}^{\prime}\in\mathcal{D}$. In such cases, the inverse deformation $F_{\text{inv-def}}$ would be unable to map the single point $\mathbf{p}^{\prime}$ back to the respective distinct points $\mathbf{p}_{1}^{\prime\prime}$ and $\mathbf{p}_{2}^{\prime\prime}$.